# Nonreciprocal tripartite entanglement and asymmetric Einstein-Podolsky-Rosen steering via directional quantum squeezing

### v1

@@ Abstract
The generation and manipulation of multipartite entanglement and EPR steering in macroscopic systems not only play a fundamental role in exploring the nature of quantum mechanics, but are also at the core of current developments of various nascent quantum technologies. Here we report a theoretical method using directional injection of quantum squeezing to produce nonreciprocal multipartite entanglement and EPR steering in a three-mode optomechanical system with closed-loop coupling. We show that by directionally applying a two-photon parametric driving field with a phase-matched squeezed vacuum reservoir to an optomechanical resonator, a squeezed optical mode can be introduced for one of its input directions, thereby yielding an asymmetric enhancement of optomechanical interaction and the time-reversal symmetry breaking of the system. Based on this feature, it is found that bipartite and tripartite entanglement and the associated EPR steering of the subsystems can only be generated when the coherent driving field input from the squeezing injection direction, namely, achieving nonreciprocity in such quantum correlations. More excitingly, it is also found that by properly adjusting the squeezing parameter, the overall asymmetry of EPR steering can be stepwise driven from no-way regime, one-way regime to two-way regime. These findings, holding promise for preparing rich types of entangled quantum resources with nonreciprocal correlations, may have potential applications in the area of quantum information processing such as quantum secure direct communication and one-way quantum computing.
@@ I Introduction
Entanglement, allowing perfectly correlated positions and momenta for two spatially separated particles, has long been intriguing in quantum physics and enables numerous advanced quantum information protocols spanning from quantum networking to quantum sensing Horodecki et al. (2009). The concept of entanglement was originally addressed by Schrödinger Schrödinger (1935) in his response to the issue of the “spooky action-at-a-distance” predicted by Einstein, Podolsky, and Rosen (EPR) in their famous paradox Einstein et al. (1935), where he also coined a term that came to be known as the EPR steering Reid (1989); Ou et al. (1992). From the perspective of violations of local-hidden-state models, Wiseman et al. has formalized an operational benchmark for EPR steering Wiseman et al. (2007), by which they further proved that under the hierarchy of quantum nonlocality, EPR steering is a strict subset of entanglement and a strict superset of Bell nonlocality. An appealing feature of EPR steering is that it describes how local measurements on one part of the system can steer (alter) the state of the other part at a different location. This defining characteristics, not held by the other two types of quantum nonlocality, reveals the intrinsic asymmetry of EPR steering and offers an insight into directional nonlocality He et al. (2015a), which plays an indispensable role in enabling quantum techniques using untrusted devices Xiang et al. (2022), such as secure quantum key distribution Branciard et al. (2012); Walk et al. (2016), randomness certification Li et al. (2024), and no-cloning quantum teleportation He et al. (2015b); Chiu et al. (2016). In the past few decades, after a series of rigorous mathematical characterizations of quantum nonlocality Simon (2000); Adesso et al. (2004); Adesso and Illuminati (2007); Kogias et al. (2015), a great deal of progresses have been made experimentally to prepare entangled or steerable states of microscopic and macroscopic particles, involving platforms based on photons Wang et al. (2016a), ions Stute et al. (2012), atoms van Leent et al. (2022), superconducting circuits Kurpiers et al. (2018), and cavity optomechanical (COM) devices Palomaki et al. (2013); Kotler et al. (2021); Mercier de Lépinay et al. (2021). However, in terms of generation and manipulation of macroscopic entanglement, it is still challenging to avoid the decoherence effect induced by device imperfection. Very recently, to overcome this obstacle and achieve entangled state with high fidelity, a large number of theoretical proposals have been raised, which relies on synthetic gauge field Lai et al. (2022a); Liu et al. (2023), reservoir engineering Wang and Clerk (2013); Yang et al. (2015); Jiao et al. , dark-mode or feedback control Lai et al. (2022b); Huang et al. (2022a); Li et al. (2017), photon counting Ho et al. (2018), dynamical modulation Wang et al. (2016b); Yang et al. (2024), and optical nonreciprocity Jiao et al. (2020, 2022); Huang et al. (2024).
In parallel, nonreciprocal optical devices, such as optical diodes and circulators, have attracted intense interests due to their potential applications in the next-generation photonic information processing Sounas and Alù (2017). The conventional approach for achieving nonreciprocal transmission of light mainly relies on magneto-optical effect Shoji and Mizumoto (2014), which requires magnetic-based materials that are bulk and quite lossy at optical frequency. In recent years, by breaking the time-reversal symmetry Caloz et al. (2018) via spatial-temporal modulation, optical nonlinearity or Sagnac effect, several magnetic-free scheme have been theoretically proposed and experimentally demonstrated based on atomic ensembles Zhang et al. (2018); Yang et al. (2019); Liang et al. (2020), optical Dong et al. (2015); Xia et al. (2018); Rosario Hamann et al. (2018) or COM Manipatruni et al. (2009); Shen et al. (2016); Bernier et al. (2017) devices, synthetic structures Peng et al. (2014); Chang et al. (2014), moving medium Wang et al. (2013); Lü et al. (2017); Maayani et al. (2018), and squeezing injectionTang et al. (2022). With these advances, the peculiar feature of nonreciprocal optical devices, allowing the flow of light from one side but block it from the other, has recently been employed to realize the unidirectional manipulation of light-matter interaction, which enables a variety of classical nonreciprocal phenomenon, such as one-way optical chaos Zhang et al. (2021) or solitons Li et al. (2021), unidirectional phonon Jiang et al. (2018) or magnon Huang et al. (2022b) lasing, and nonreciprocal enhancement optomechanical sidebands Li et al. (2020), to name a few. Very recently, this functionality has also been extended into the quantum regime to achieve nonreciprocal control of diverse nonclassical effect, such as photon Huang et al. (2018); Yang et al. (2023); Wang et al. (2023) or magnon Wang et al. (2022) blockade, microscopic Orr et al. (2023); Shen et al. (2023); Bin et al. (2024) or macroscopic Jiao et al. (2020, 2022); Huang et al. (2024) entanglement, quantum phase transition Fruchart et al. (2021); Zhu et al. (2024), and EPR steering Guan et al. (2024); Wu et al. (2024); Zhong et al. (2023).
Inspired by these studies, we investigate how to achieve nonreciprocal entanglement and asymmetric EPR steering in a hybrid COM system based on whispering-gallery-mode (WGM) resonator. We show that by applying a two-photon parametric driving field from one input direction while not applying it from the other, the two degenerate counterpropagating optical modes of the WGM resonator can be unidirectionally squeezed, leading to the time-reversal symmetry breaking of the system. Based on this chiral squeezing characteristics, we show that various types of bipartite entanglement of the subsystems and the tripartite entanglement of the whole system can be generated only when coherent light input from the squeezing injection direction, which indicates the achievement of nonreciprocal entanglement. More interestingly, by properly adjusting the squeezing parameters, we further show that the directionality of EPR steering can also be stepwise driven from no-way regime, one-way regime to two-way regime in an asymmetric way. These results, opening up a promising way to prepare entangled quantum resources with nonreciprocal and asymmetric features, are useful for a variety of secure quantum information protocols Branciard et al. (2012); Walk et al. (2016); Li et al. (2024); He et al. (2015b); Chiu et al. (2016).
@@
This paper is structured as follows. In Sec. II, we introduce the theoretical model of the proposed COM system and obtain the effective Hamiltonian and master equation of this system, by which we calculate the system dynamics and evaluate the quantitative measures for entanglement and EPR steering. In Sec. III, based on the numerical simulations, we analyze the method and mechanism for achieving nonreciprocal generation and manipulation of various types of entanglement and EPR steering. In Sec. IV, a brief summary is given. The detailed derivation process of the effective Hamiltonian, the associated master equation, and the system dynamics is given in the Appendix.
Figure 1: Schematic of the proposed COM setup consisting of a bottle WGM resonator, an optical waveguide, and a tapered glass-fiber nanospike. The WGM resonator supports two degenerate counterpropagating optical modes, i.e., CW and CCW modes. The optical driving field is input from the waveguide, which can be coupled into and out of the WGM resonator via evanescent coupling. When placing a nanospike close to the WGM resonator, it excites two mechanical modes due to the COM interaction induced by the WGM evanescent field, whose vibrating directions are parallel and orthogonal to the resonator surface, respectively. To achieve nonreciprocal manipulation of quantum entanglement and EPR steering, we consider two different scenarios for the the CW and CCW input cases. For the CW input case, only a coherent driving field is applied, while for the CCW input case, an additional two-photon parametric driving field with a phase-matched squeezed vacuum reservoir is applied.
@@ II Theoretical model and system dynamics
The proposed setup consists of a bottle WGM resonator, an optical waveguide, and a tapered glass-fiber nanospike, which is depicted in Fig. 1. Due to the spatial symmetry of the bottle WGM resonator, it can support two degenerate counterpropagating optical modes with resonance frequency $\omega_{c}$, i.e., the clockwise (CW) and counterclockwise (CCW) modes, respectively. The nanospike is usually fabricated by scanning an oxybutane flame along the length of a glass fiber while gently pulling it. As demonstrated in the recent experiments Xie et al. (2016, 2019); Pennetta et al. (2020), by mounting such nanospike on a stainless-steel holder and placing it close to the bottle WGM resonator, it offers two mechanical modes due to the WGM evanescent field induced COM interaction. These two mechanical modes with frequency $\omega_{m,j}$ $(j=1,2)$ are vibrating in the directions that are orthogonal and parallel to the bottle surface, and they are coupled linearly due to the optically mediated hybridization effect Shkarin et al. (2014); Lin et al. (2010). The WGM resonator can be driven from both sides, where the optical driving fields from the waveguide are coupled into and out of the WGM resonator via evanescent coupling.
@@
In addition, we consider two different scenarios for the driving fields applied from the CW and CCW input directions, respectively. For the CCW input case, a coherent driving field with frequency $\omega_{d}/2$ and a two-photon parametric driving field with frequency $\omega_{d}$ are simultaneously applied to the waveguide. In this situation, in a rotating reference frame with respect to the driving frequency $\omega_{d}/2$, the Hamiltonian of the whole system can be expressed as (setting $\hbar=1$)
$\hat{H}_{\circlearrowleft}=$ $\hat{H}_{\textrm{sys}}+\hat{H}_{\textrm{dr},\circlearrowleft},$
$\hat{H}_{\textrm{sys}}=$ $\Delta_{c}\hat{a}_{\circlearrowleft}^{\dagger}\hat{a}_{\circlearrowleft}+\sum_{j=1,2}\left[\dfrac{\omega_{m,j}}{2}(\hat{p}_{j}^{2}+\hat{q}_{j}^{2})-g_{j}\hat{a}_{\circlearrowleft}^{\dagger}\hat{a}_{\circlearrowleft}\hat{q}_{j}\right]$
$+\chi\hat{q}_{1}\hat{q}_{2},$
$\hat{H}_{\textrm{dr},\circlearrowleft}=$ $\Xi_{d}(e^{-i\theta_{d}}\hat{a}_{\circlearrowleft}^{\dagger 2}+e^{i\theta_{d}}\hat{a}_{\circlearrowleft}^{2})+i\varepsilon_{d,\circlearrowleft}(\hat{a}_{\circlearrowleft}^{\dagger}-\hat{a}_{\circlearrowleft}),$ (1)
where $\hat{a}_{\circlearrowleft}$ ($\hat{a}_{\circlearrowleft}^{\dagger}$) is the annihilation (creation) operator of the CCW mode, while $\hat{q}_{j}$ and $\hat{p}_{j}$ are the dimensionless displacement and momentum operators of the $j$th mechanical mode, respectively. $\Delta_{c}=\omega_{c}-\omega_{d}/2$ denotes the optical detuning, and $\omega_{c}$ is the resonance frequency of the CCW (CW) mode. $g_{j}=(\omega_{c}/R)\sqrt{\hbar/m_{j}\omega_{m,j}}$ is the single-photon COM coupling strength, with $R$ the bottle radius and $m_{j}$ the effective mass of the $j$th mechanical mode. $\chi$ is the mechanical coupling strength. $\Xi_{d}$ and $\theta_{d}$ are the amplitude and phase of the two-photon parametric driving field. The amplitude of the coherent driving field is given by $|\varepsilon_{d,\circlearrowleft}|=\sqrt{2\kappa P_{d}/\hbar\omega_{d}}$, where $P_{d}$ is the input laser power and $\kappa$ is the optical decay rate.
@@
After performing the Bogoliubov transformation with a unitary operator, $S(\eta_{d})=\exp[(-\eta_{d}\hat{a}_{\circlearrowleft}^{\dagger 2}+\eta_{d}^{\ast}\hat{a}_{\circlearrowleft}^{2})/2]$, a squeezed optical mode $\hat{a}_{s,\circlearrowleft}$ can be introduced, i.e.,
$S^{\dagger}(\eta_{d})\hat{a}_{\circlearrowleft}S(\eta_{d})=\cosh(r_{d})\hat{a}_{s,\circlearrowleft}-e^{-i\theta_{d}}\sinh(r_{d})\hat{a}_{s,\circlearrowleft}^{\dagger},$ (2)
@@
where $\eta_{d}=r_{d}e^{-i\theta_{d}}$ is the complex squeezing parameter, with a squeezing strength $r_{d}$ and a squeezing reference angle $\theta_{d}$. Hence, by dropping the constant terms, the effective Hamiltonian of the total system in the squeezing picture is derived as [see Appendix A for more details]
@@
$\left.+\dfrac{\zeta_{p,j}}{2}(e^{-i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{\dagger 2}+e^{i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{2})\hat{q}_{j}-F_{j}\hat{q}_{j}\right]+\chi\hat{q}_{1}\hat{q}_{2}$
@@
where $\omega_{s}=(\Delta_{c}-2\Xi_{d})\exp(2r_{d})$ is the effective resonance frequency of the squeezed optical mode. $\zeta_{s,j}=g_{j}\cosh(2r_{d})$ and $\zeta_{p,j}=g_{j}\sinh(2r_{d})$ are the effective COM coupling strength for the $j$th mechanical mode induced by the radiation pressure and the parametric amplification, respectively. $F_{j}=g_{j}\sinh^{2}(r_{d})$ denotes the strength of the constant mechanical driving force exerted on the $j$th mechanical mode, which is caused by the parametric amplification.
As discussed in the preceding studies Jiao et al. ; Tang et al. (2022); Guan et al. (2024); Wu et al. (2024); Lü et al. (2015), by adjusting the squeezing parameters with respect to the two-photon parametric driving field, the effective light-matter interaction can be manipulated and enhanced in a controllable way. However, it is worth emphasizing that when the squeezed optical mode is coupled to a normal vacuum reservoir, the optical dissipation will be inevitably amplified in the meantime, thus leading to its quantum decoherence. To eliminate such optical dissipation induced by the intracavity squeezing, one can apply a broadband squeezed-vacuum reservoir that is phase-matched with the squeezed optical mode, i.e., the squeezing strength $r_{e}$ and phase $\theta_{e}$ of this reservoir satisfying the relationships $r_{e}-r_{d}=0$ and $\theta_{e}-\theta_{d}=\pm n\pi~{}(n=1,3,5,\ldots)$. Note that the injection of this required broadband squeezed-vacuum reservoir has been reported through using an optical parametric amplifier Kashiwazaki et al. (2020). Additionally, given that the two mechanical modes are coupled with two independent thermal reservoirs at same bath temperature $T$, the Born-Markovian master equation describing the dynamics of the whole system in the squeezing picture can be derived as [see Appendix B for more details]

### v2

# Phase-selective tripartite entanglement and asymmetric Einstein-Podolsky-Rosen steering in squeezed optomechanics

@@ Abstract
The generation and manipulation of multipartite entanglement and EPR steering in macroscopic systems not only play a fundamental role in exploring the nature of quantum mechanics, but are also at the core of current developments of various nascent quantum technologies. Here we report a theoretical method using squeezing-phase-controlled quantum noise flows to selectively generate and manipulate quantum entanglement and asymmetric EPR steering in a nonlinear $\chi^{(2)}$ whispering-gallery-mode (WGM) optomechanical resonator. We show that by pumping the $\chi^{(2)}$ nonlinear medium with two-photon optical fields and broadband squeezed lights, a pair of counterpropagating squeezed optical modes could be introduced to the WGM resonator, each coupled with an independent squeezed vacuum reservoir. This configuration could enable squeezing-phase-controlled light-reservoir interaction for each squeezed optical mode, providing a flexible tool for tailoring asymmetric optical noise flows in the counterpropagating modes. Based on this unique feature, it is found that with the injection of asymmetric noise flows, the generation of various types of bipartite and tripartite entanglement become phase-dependent and thus they can be produced in an asymmetric way. More excitingly, it is also found that by further properly adjusting the squeezing parameters, the overall asymmetry of EPR steering can also be stepwise driven from no-way regime, one-way regime to two-way regime. These findings, holding promise for preparing rich types of entangled quantum resources with asymmetric features, may have potential applications in the area of secure quantum information processing such as quantum secure direct communication and one-way quantum computing.
@@ I Introduction
Entanglement, allowing perfectly correlated positions and momenta for two spatially separated particles, has long been intriguing in quantum physics and enables numerous advanced quantum information protocols spanning from quantum networking to quantum sensing [1]. The concept of entanglement was originally addressed by Schrödinger [2] in his response to the issue of the “spooky action-at-a-distance” predicted by Einstein, Podolsky, and Rosen (EPR) in their famous paradox [3], where he also coined a term that came to be known as the EPR steering [4, 5]. From the perspective of violations of local-hidden-state models, Wiseman et al. has formalized an operational benchmark for EPR steering [6], by which they further proved that under the hierarchy of quantum nonlocality, EPR steering is a strict subset of entanglement and a strict superset of Bell nonlocality. An appealing feature of EPR steering is that it describes how local measurements on one part of the system can steer (alter) the state of the other part at a different location. This defining characteristics, not held by the other two types of quantum nonlocality, reveals the intrinsic asymmetry of EPR steering and offers an insight into directional nonlocality [7], which plays an indispensable role in enabling quantum techniques using untrusted devices [8], such as secure quantum key distribution [9, 10], randomness certification [11], and no-cloning quantum teleportation [12, 13]. In the past few decades, after a series of rigorous mathematical characterizations of quantum nonlocality [14, 15, 16, 17], a great deal of progresses have been made experimentally to prepare entangled or steerable states of microscopic and macroscopic particles, involving platforms based on photons [18], ions [19], atoms [20], superconducting circuits [21], and cavity optomechanical (COM) devices [22, 23, 24]. However, in terms of generation and manipulation of macroscopic entanglement, it is still challenging to avoid the decoherence effect induced by device imperfection. Very recently, to overcome this obstacle and achieve entangled state with high fidelity, a large number of theoretical proposals have been raised, which relies on synthetic gauge field [25, 26], reservoir engineering [27, 28], dark-mode or feedback control [29, 30, 31], photon counting [32], dynamical modulation [33, 34], injection of quantum squeezing [35, 36], and optical nonreciprocity [37, 38, 39].
On the other hand, squeezed light, characterized by reduced quantum noise in selected quadratures, has emerged as a vital resource in modern quantum science [40]. Squeezed light has been widely utilized to cool mechanical motion [41], enhance light-matter interaction [42, 43, 44], and advance quantum information processing [45, 46] across a broad range of quantum platforms, including optomechanical devices [47], cavity QED systems [48], and trapped ions [49]. In particular, the peculiar feature of quadrature noise reduction make squeezed light an indispensable resource for improving the sensitivity of quantum measurements, leading to practical applications in gravitational-wave detection [50], nanoscale sensing [51], and nondemolition qubit readout [52]. In practice, squeezed light is typically introduced to quantum systems either by external injection or by internal generation via $\chi^{(2)}$ nonlinearity. However, for current experimental techniques, both approaches face significant limitations: external injection suffers from inevitable transmission and coupling losses that render squeezed states fragile, while intracavity generation is fundamentally constrained by the 3 dB limit [53], which prevents quantum noise of a cavity field from being reduced below half of the zero-point fluctuations. Recently, Lü et al. proposed a novel squeezed-light scheme that combines external injection and intracavity generation [54], effectively addressing the limitations of the conventional methods and enabling single-photon quantum processes in the weak-coupling regime of an optomechanical system. Building on this approach, subsequent studies have further explored the potential of this hybrid squeezing strategy in various quantum applications, such as surpassing the 3 dB limit of intracavity squeezing [55], enhancing qubit readout [56], and generating long-lived cat states [57, 58].
Inspired by these studies, we here investigate how to achieve selective generation of tripartite entanglement and asymmetric EPR steering through tuning squeezing phase in a WGM optomechanical resonator. The WGM resonator is made of $\chi^{(2)}$ nonlinear materials and supports two degenerate counterpropagating optical modes and two mechanical modes. Particularly, we show that by pumping the $\chi^{(2)}$ nonlinear medium with two-photon optical fields and broadband squeezed lights, a pair of counterpropagating squeezed optical modes could be introduced to the WGM resonator, each coupled with an independent squeezed vacuum reservoir. This configuration enables a squeezing-phase-controlled light-reservoir interaction for each squeezed optical mode and provides a flexible tool for manipulating the flow of optical input noises,which plays a key role in producing phase-selective entanglement. Particularly, it is found that when tailoring asymmetric flow of optical noises in different input directions, the generation of various types of bipartite and tripartite entanglement become phase-dependent and they can only be generated in an asymmetric way. More interestingly, by properly adjusting the squeezing phase-matched condition, we further show that the overall asymmetry of EPR steering can be stepwise driven from no-way regime, one-way regime to two-way regime. These generated asymmetric entangled states, with their rich variety, are applicable to a wide range of on-chip secure quantum information protocols [9, 10, 11, 12, 13]. Besides, our work, providing an efficient all-optical approach to engineer quantum states in an asymmetric way, can also be extended to explore the manipulation of other types of quantum effect, such as photon blockade [59, 60, 61] and quantum phase transition [62, 63]. As such, we believe that our proposal would not only be promising for serving in fundamental tests of quantum theories, but also can serve as key quantum resources for nascent quantum technologies like quantum secure communication [12, 13] and one-way quantum computing [64, 65].
Figure 1: (a) Schematic of a squeezed COM setup, showing a tapered-fibers-interfaced microdisk WGM resonator dispersively coupled to a string of nanomechanical oscillators. The WGM resonator is made of $\chi^{(2)}$ nonlinear materials, supporting two degenerate counterpropagating modes. The mechanical string offers two mechanical modes, which can be excited by placing the string in the evanescent field of the WGM resonator. When applying two-photon pumping fields along with broadband squeezed lights from the right fiber ports, two squeezed counterpropagating modes $\hat{a}_{s,\circlearrowright}$ and $\hat{a}_{s,\circlearrowleft}$ could be introduced in the WGM resonator, which are coupled with independent squeezed vacuum reservoirs. The coherent driving field is input from the left fiber port and is coupled into or out of the WGM resonator via evanescent coupling. (b) Density plot of the enhancement factor $\Pi$ as a function of squeezing strength $r_{d}$ and squeezing reference angle $\theta_{d}$. The effective COM interaction can be either enhanced ($\Pi>1$) or reduced ($\Pi<1$) depending on squeezing parameters. (c) Plot of the effective optical thermal noise $N_{s,i}$ and the effective two-photon correlation strength $M_{s,i}$ versus the phase difference $\delta\theta_{i}$ for different values of $\delta r_{i}$. Specifically, for the CCW input case, we consider the squeezed optical mode $\hat{a}_{s,\circlearrowleft}$ is phase-matched with its squeezed vacuum reservoir, leading to $N_{s,\circlearrowleft}=M_{s,\circlearrowleft}=0$. For the CW input case, it is phase-mismatched for $\hat{a}_{s,\circlearrowright}$, leading to $N_{s,\circlearrowright}\neq 0$ and $M_{s,\circlearrowright}\neq 0$. (d) The steady-state CM $V$ for opposite input cases. The parameters used here are given in the main text.
@@
This paper is structured as follows. In Sec. II, we introduce the theoretical model of the proposed COM system and obtain the effective Hamiltonian and master equation of this system, by which we calculate the system dynamics and evaluate the quantitative measures for entanglement and EPR steering. In Sec. III, based on the numerical simulations, we analyze the method and mechanism for achieving phase-selective generation and manipulation of various types of entanglement and EPR steering. In Sec. IV, a brief summary is given. The detailed derivation process of the effective Hamiltonian, the associated master equation, and the system dynamics is given in the Appendix.
@@ II Theoretical model and system dynamics
In this paper, we propose how to use phase-controlled quantum noise flows to selectively generate and manipulate quantum entanglement and asymmetric EPR steering in a three-mode COM system with two mechanical modes. Here, as depicted in Fig. 1(a), we consider a near-field COM system as a possible experimental setup [66, 67], which consists of a tapered-fibers-interfaced microdisk WGM resonator dispersively coupled to a string of nanomechanical oscillators. The microdisk WGM resonator is assumed to be made up of $\chi^{(2)}$ nonlinear materials, e.g., lithium niobate or aluminum nitride, thereby supporting optical parametric down-conversion process. Due to the spatial symmetry of the microdisk WGM resonator, it supports two degenerate counterpropagating optical modes with resonance frequency $\omega_{c}$, i.e., the clockwise (CW) and counterclockwise (CCW) modes, respectively. For high-quality WGM resonators without backscattering losses, the CW and CCW modes could keep uncoupled. The two mechanical modes with fundamental frequency $\omega_{m,j}$ $(j=1,2)$ are offered by the nanomechanical oscillators in the mechanical string. As demonstrated in the experiment [66], by placing such mechanical string close to the tightly confined evanescent field of the microdisk WGM resonator, each oscillator can excite a mechanical vibration mode in the direction orthogonal to the microdisk surface, which is resulting from the evanescent field induced COM interaction. Furthermore, due to the high-finesse of optical and mechanical modes, as well as the mitigated losses and thermal effects during the coupling process, the observed COM interaction in this experiment is purely dispersive. In addition, given that the two mechanical modes are degenerate and interact with the same cavity mode, an effective phonon-hopping interaction could be generated between them, which is mediated by the cavity field induced hybridization effect [68]. Note that such optomechanically induced hybridization of distinct mechanical modes has been experimentally demonstrated in nano-optomechanical structures [69], WGM resonators [70, 71], and multimode circuit systems [72].
@@
As shown in Fig. 1(a), the microdisk WGM resonator can be driven from two opposite input directions. The coherent driving field with frequency $\omega_{d}/2$ applied to port $1$ (port $2$) could be evanescently coupled into the CCW (CW) mode, which is referred to as the CCW (CW) input case. For the CCW input case, the Hamiltonian of the whole system in a frame rotating with $\omega_{d}/2$ can be expressed as (setting $\hbar=1$)
$\hat{H}_{\circlearrowleft}=$ $\hat{H}_{\textrm{c}}+\hat{H}_{\textrm{m}}+\hat{H}_{\textrm{om}}+\hat{H}_{\textrm{dr},\circlearrowleft},$
$\hat{H}_{\textrm{c}}=$ $\Delta_{c}\hat{a}_{\circlearrowleft}^{\dagger}\hat{a}_{\circlearrowleft}+\Xi_{d}(e^{-i\theta_{d}}\hat{a}_{\circlearrowleft}^{\dagger 2}+e^{i\theta_{d}}\hat{a}_{\circlearrowleft}^{2}),$
$\hat{H}_{\textrm{m}}=$ $\dfrac{\omega_{m,1}}{2}(\hat{p}_{1}^{2}+\hat{q}_{1}^{2})+\dfrac{\omega_{m,2}}{2}(\hat{p}_{2}^{2}+\hat{q}_{2}^{2})+\lambda\hat{q}_{1}\hat{q}_{2},$
$\hat{H}_{\textrm{om}}=$ $-g_{1}\hat{a}_{\circlearrowleft}^{\dagger}\hat{a}_{\circlearrowleft}\hat{q}_{1}-g_{2}\hat{a}_{\circlearrowleft}^{\dagger}\hat{a}_{\circlearrowleft}\hat{q}_{2},$
$\hat{H}_{\textrm{dr},\circlearrowleft}=$ $i\varepsilon_{d,\circlearrowleft}(\hat{a}_{\circlearrowleft}^{\dagger}-\hat{a}_{\circlearrowleft}),$ (1)
[…]
In our proposed scheme, the primary distinction between the CW and CCW input cases lies in the squeezing phase and strength of their injected squeezed vacuum reservoirs, which could induce asymmetric quantum noise flows in these counterpropagating modes. In the following, we discuss how to independently inject squeezed vacuum reservoirs to CW and CCW modes, and then derive the effective master equation for CCW input case as a specific example. As shown in Fig. 1(a), we consider applying two different broadband squeezed optical fields to these counterpropagating modes, which are centred around frequency $\omega_{c}$ and characterized by squeezing strengths $r_{e,i}$ and reference phase angles $\theta_{e,i}$ ($i=\circlearrowleft,\circlearrowleft$), respectively. In practice, these broadband squeezed optical fields can act as two independent squeezed-vacuum reservoirs to CW and CCW modes [74]. This is because the bandwidth of squeezed optical field can be up to gigahertz (GHz[48], while the typical linewidth of WGM resonators is of the order of megahertz (MHz). In this case, the broadband squeezed optical field can be well approximated as having infinite bandwidth to WGM optical modes and be safely treated as a squeezed vacuum reservoir. Besides, the two mechanical modes are assumed to be coupled with two independent thermal reservoirs at the same bath temperature $T$. Then, in terms of the CCW input case, including the dissipations of the optical and mechanical modes, the dynamics of the total system is governed by the following Born-Markovian master equation [75]
$\dot{\rho}_{\circlearrowleft}=$ $i\big{[}\rho_{\circlearrowleft},\hat{H}_{\circlearrowleft}\big{]}+\frac{\kappa}{2}(N_{e,\circlearrowleft}+1)\mathcal{D}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}+\frac{\kappa}{2}N_{e,\circlearrowleft}\mathcal{D}[\hat{a}_{\circlearrowleft}^{\dagger}]\rho_{\circlearrowleft}$
$-\frac{\kappa}{2}M_{e,\circlearrowleft}\mathcal{G}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}-\frac{\kappa}{2}M_{e,\circlearrowleft}^{\ast}\mathcal{G}[\hat{a}_{\circlearrowleft}^{\dagger}]\rho_{\circlearrowleft}$
where
$\mathcal{G}[\hat{a}_{\circlearrowleft}]\rho_{\circlearrowleft}$ $=2\hat{a}_{\circlearrowleft}\rho_{\circlearrowleft}\hat{a}_{\circlearrowleft}-(\hat{a}_{\circlearrowleft}\hat{a}_{\circlearrowleft}\rho_{\circlearrowleft}+\rho_{\circlearrowleft}\hat{a}_{\circlearrowleft}\hat{a}_{\circlearrowleft})$ (3)
[…]
@@
After performing the Bogoliubov transformation with a unitary operator, $U_{\circlearrowleft}(\eta_{d})=\exp[(-\eta_{d}\hat{a}_{\circlearrowleft}^{\dagger 2}+\eta_{d}^{\ast}\hat{a}_{\circlearrowleft}^{2})/2]$, a squeezed optical mode $\hat{a}_{s,\circlearrowleft}$ can be introduced, i.e.,
$U_{\circlearrowleft}^{\dagger}(\eta_{d})\hat{a}_{\circlearrowleft}U_{\circlearrowleft}(\eta_{d})=\cosh(r_{d})\hat{a}_{s,\circlearrowleft}-e^{-i\theta_{d}}\sinh(r_{d})\hat{a}_{s,\circlearrowleft}^{\dagger},$ (4)
@@
where $\eta_{d}=r_{d}e^{-i\theta_{d}}$ is the complex squeezing parameter, with a squeezing strength $r_{d}=(1/4)\ln[(\Delta_{c}+2\Xi_{d})/(\Delta_{c}-2\Xi_{d})]$ and a squeezing reference angle $\theta_{d}$. Hence, by dropping the constant terms, the effective Hamiltonian with respect to the CCW input case in the squeezing picture is derived as [see Appendix A for more details]
@@
$\left.+\dfrac{\zeta_{p,j}}{2}(e^{-i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{\dagger 2}+e^{i\theta_{d}}\hat{a}_{s,\circlearrowleft}^{2})\hat{q}_{j}-F_{j}\hat{q}_{j}\right]+\lambda\hat{q}_{1}\hat{q}_{2}$
@@
where $\omega_{s}=(\Delta_{c}-2\Xi_{d})\exp(2r_{d})$ is the effective resonance frequency of the squeezed optical mode. $\zeta_{s,j}=g_{j}\cosh(2r_{d})$ and $\zeta_{p,j}=g_{j}\sinh(2r_{d})$ are the effective COM coupling strength for the $j$th mechanical mode induced by the radiation pressure and the parametric amplification, respectively. $F_{j}=g_{j}\sinh^{2}(r_{d})$ denotes the strength of the constant mechanical driving force exerted on the $j$th mechanical mode, which is caused by the parametric amplification. Correspondingly, the effective master equation in the squeezing picture is derived as [see Appendix B for more details]
$\dot{\tilde{\rho}}_{\circlearrowleft}=$ $U^{\dagger}(\eta_{d})\dot{\rho}_{\circlearrowleft}U(\eta_{d})$
$=$ $i\big{[}\tilde{\rho}_{\circlearrowleft},\hat{\tilde{H}}_{\circlearrowleft}\big{]}\!+\!\frac{\kappa}{2}(N_{s,\circlearrowleft}+1)\mathcal{D}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}\!+\!\frac{\kappa}{2}N_{s,\circlearrowleft}\mathcal{D}[\hat{a}_{s,\circlearrowleft}^{\dagger}]\tilde{\rho}_{\circlearrowleft}$
$-\frac{\kappa}{2}M_{s,\circlearrowleft}\mathcal{G}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}-\frac{\kappa}{2}M_{s,\circlearrowleft}^{\ast}\mathcal{G}[\hat{a}_{s,\circlearrowleft}]\tilde{\rho}_{\circlearrowleft}$
@@
$N_{s,\circlearrowleft}=$ $\sinh^{2}(r_{d})\cosh^{2}(r_{e,\circlearrowleft})+\cosh^{2}(r_{d})\sinh^{2}(r_{e,\circlearrowleft})$
$+\dfrac{1}{2}\cos\delta\theta\sinh(2r_{d})\sinh(2r_{e,\circlearrowleft}),$
$M_{s,\circlearrowleft}\!=$ $e^{i\theta_{d}}\![\sinh(r_{d})\!\cosh(r_{e,\circlearrowleft})\!+\!e^{i\delta\theta_{\circlearrowleft}}\!\cosh(r_{d})\!\sinh(r_{e,\circlearrowleft})\!]$
×$\times\![\cosh(r_{d})\!\cosh(r_{e,\circlearrowleft})\!+\!e^{-i\delta\theta_{\circlearrowleft}}\!\sinh(r_{d})\!\sinh(r_{e,\circlearrowleft})],$ (7)